Explaining Expressions for Arithmetic arithmetic sequence whose first entry is 5 and s CA-208 Chapter 9 Algebra Class Activity 9BB Sequences increases by 3 Entry Number Entry Ist 2nd 3rd 4th 5th Nth I1 14 17 a. If there were a Oth entry, what would it be? Put it in the table. b. Fill in the blanks to describe how to get entries in the sequence by starting from the Oth entr To find the Ist entry: Start at and add1 time. . To find the 2nd entry: Start at-, and add- -2 times. To find the 3rd entry: Start at--and add--3 times. . To find the 4th entry: Start at-and add--4 times. . To find the 5th entry: Start at-_ and add--5 times. . To find the Mth entry: Start at _ and add--N times. c. For each bullet in part (b), write an expression (using addition and multiplication) that corresponds to the description for finding the entry in the sequence Ist entry = 2nd entry = 3rd entry = 4th entry = 5th entry Nth entry
media%2F2cf%2F2cf582d8-b8f6-43bf-94dc-06
Answer #1

1. From table if an denote nth term of sequence then

a_1 = 5 , a_2 = 8 , a_3 = 11 , a_4 = 14 , a_5 = 17

(a) Common difference of given sequence is

8 - 5 = 11 - 8 = 14 - 11 = 17 - 14 = 3

So 0 th entry of sequence is:

a_0 = 5 - 3 = 2

(b) To find 1 st entry start at 2 and add 3 , 1 time.

To find 2 nd entry start at 2 and add 3 , 2 time.

To find 3 rd entry start at 2 and add 3 , 3 time.

To find 4 th entry start at 2 and add 3 , 4 time.

To find 5 th entry start at 2 and add 3 , 5 time.

To find n th entry start at 2 and add 3 , n time.

(c) 1 st entry = 2+ 3*1

2 nd entry = 2 + 3*2

3rd entry = 2 + 3*3

4th entry = 2 + 3*4

5th entry = 2 + 3*5

n th entry = 2 + 3*n

2.given sequence is 2. 5 , 8 , 11 , 14 , .................................

Difference between each consecutive term is 5-2 = 8-5 = 11 - 8 = 14 - 11 = 3

So difference between each consecutive term is 3 which is same. So given sequece represents Airthmetic series.

General term of any airthmetic series is

a_n = a+ (n-1)d

In given series a = 2 , d = 3

So general term is:

a_n = 2+ (n-1)3

= 2+3n - 3

  = 3n - 1